Find the inverse Laplace of {2s+5/s^2+25}?

A) 2Sin5t+Cos5t

B) Cos5t-2Sin5t

C) 2Cos5t+Sin5t

D) 2Cos5t-Sin5t

1
Expert's answer
2022-02-11T10:11:45-0500

Let us find the inverse Laplace of "\frac{2s+5}{s^2+25}=2\cdot\frac{s}{s^2+25}+\frac{5}{s^2+25}." Taking into account that "\frac{s}{s^2+b^2}\to\cos (bt),\" and "\frac{b}{s^2+b^2}\to\sin (bt)," we conclude that the inverse Laplace of "\frac{2s+5}{s^2+25}" is

"2\cos(5t)+\sin(5t)."


Answer: C)